OpenStudy (anonymous):

Dylan's Bicycle Shop offers a new bicycle on sale for $259.92 after a 10% markdown. What was the cost of the bicycle to Dylan if the markup was 30% based on the cost? (Round to the nearest cent.)

OpenStudy (anonymous):

anbody no this

OpenStudy (anonymous):

259.92 = Original Price * 0.9 259.92 = (1.3 * Cost to Dylan) * 0.9 259.92 = 1.3 * 0.9 * x x = cost to Dylan x = 259.92 / (1.3*0.9) = $222.15

OpenStudy (anonymous):

thank you